sqrt(x+3)+ sqrt(2x-3)=6
Answer by AnlytcPhil(1810)


 ___    ____
Öx+3 + Ö2x-3 = 6

Isolate either radical term
        ____        ___ 
       Ö2x-3 = 6 - Öx+3

Square both sides:
        ____           ___ 
      (Ö2x-3)² = (6 - Öx+3)²

It's easy to square the left side, for the squaring
just cancels the square root.  However it's not so
easy to square the right side.  Put it down twice
and use FOIL
                       ___       ___
          2x-3 = (6 - Öx+3)(6 - Öx+3)

On the right:

"F" = 6·6 = 36
            ___       ___
"O" = (6)(-Öx+3) = -6Öx+3
         ___          ___
"I" = (-Öx+3)(6) = -6Öx+3
         ___    ___       ___
"L" = (-Öx+3)(-Öx+3) = (-Öx+3)² = x+3 

So we have:
                       ___     ___
         2x-3 = 36 - 6Öx+3 - 6Öx+3 + x+3
                        ___
         2x-3 = 36 - 12Öx+3 + x+3
                        ___
         2x-3 = 39 - 12Öx+3 + x

Isolate the radical term
          ___
       12Öx+3 = 39 + x - 2x + 3
          ___
       12Öx+3 = 42 - x 

Square both sides:
         ___
     (12Öx+3)² = (42 - x)²

      144(x+3) = (42 - x)(42 - x)

    144x + 432 = 1764 - 42x - 42x + x²

    144x + 432 = 1764 - 84x + x²

Get 0 on the left:

             0 = x² - 228x + 1332

The right side may factor. However since the
numbers are so big, it's probably easier to use
the quadratic formula:
                __________________
     -(-228) ± Ö(-228)²-4(1)(1332)
x = -------------------------------
                2(1)
             _____
      228 ± Ö46656
x = ----------------
            2
             
      228 ± 216
x = -------------
          2
              
Using the +

      228 + 216
x = -------------
          2

     444
x = ----- = 222
      2

Using the -

      228 - 216
x = -------------
          2

     12
x = ---- = 6
      2

Now we must check, because often we get extraneous 
answers in equations with even roots. Checking the
answer x = 222

           ___    ____
          Öx+3 + Ö2x-3 = 6
     _____    ________
    Ö222+3 + Ö2(222)-3 = 6
            ___    ___
           Ö225 + Ö441 = 6

               15 + 21 = 6
 
                    36 = 6

No that doesn't check. So we discard that answer.
Checking the answer x=6

           ___    ____
          Öx+3 + Ö2x-3 = 6
         ___    ______
        Ö6+3 + Ö2(6)-3 = 6
                _    _
               Ö9 + Ö9 = 6

                 3 + 3 = 6
 
                     6 = 6

That checks.  So there is one solution,
 
x = 6

Edwin